OpenStudy (kellyspeakslouder):

f(x) = 4x + 7, g(x) = 3x2 Find (f + g)(x). A. 3x2 + 4x + 7 B. -3x2 + 4x + 7 C. Quantity four x plus 7 divided by three x squared. D. 12x3 + 21x

OpenStudy (kellyspeakslouder):

@zepdrix

OpenStudy (kellyspeakslouder):

@pooja195

OpenStudy (kellyspeakslouder):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

(f+g)(x) is the same as f(x) + g(x)

jimthompson5910 (jim_thompson5910):

so you just add the two right sides of the functions

jimthompson5910 (jim_thompson5910):

Then arrange the terms

OpenStudy (kellyspeakslouder):

...A? I'm not sure.. I am horrible at math.

jimthompson5910 (jim_thompson5910):

Correct

OpenStudy (kellyspeakslouder):

really??

jimthompson5910 (jim_thompson5910):

\[\Large {\color{red}{f(x) = 4x+7}}\] \[\Large {\color{blue}{g(x) = 3x^2}}\] ------------------------------------------ \[\Large (f+g)(x) = f(x) + g(x)\] \[\Large (f+g)(x) = {\color{red}{f(x)}} + {\color{blue}{g(x)}}\] \[\Large (f+g)(x) = {\color{red}{4x+7}} + {\color{blue}{3x^2}}\] \[\Large (f+g)(x) = {\color{blue}{3x^2}}+{\color{red}{4x+7}}\] \[\Large (f+g)(x) = 3x^2+4x+7\]

OpenStudy (kellyspeakslouder):

thank you so much!

jimthompson5910 (jim_thompson5910):

no problem